OpenStudy (anonymous):

Which of the following is a solution to square root of x minus 2 end root plus 5 equals 1 ?

OpenStudy (anonymous):

\[\sqrt{x-2}+5=1\] ????

OpenStudy (anonymous):

\[\sqrt{x-2}=1-5\] \[\sqrt{x-2}=-4\] by suqaring both sides \[(\sqrt{x-2})^2=(-4)^2\] x-2=16 x=18